OpenStudy (anonymous):

Lannette drove 248 miles in 4 hours. At that rate, how long will it take her to drive an additional 93 miles? a. 2 hours c. 3 hours b. 1.5 hours d. 2.5 hours

OpenStudy (anonymous):

i just need the problem set up for me to do it.

OpenStudy (anonymous):

Set up a proportion: miles driven/time = miles driven/time 248/4 = 93/t where t = time in hours Now solve for t.

OpenStudy (anonymous):

ohkay. t=1.5

OpenStudy (anonymous):

You're welcome.